# Advection selects pattern in multi-stable emulsions of active droplets

### v1

@@ Abstract
The relative importance of advection and diffusion can be quantified by the viscous and diffusive dissipation, respectively. Fig. 4D–F show that advection is relevant early, but decays with time, essentially independent of $k$ (blue lines are similar in panels D–F). The diffusive dissipation also peaks early, and decays similarly to the viscous dissipation during droplet coarsening. However, when reactions are present, the diffusive dissipation plateaus at a much larger value compared to the viscous dissipation, indicating that energy induced by reactions leads to diffusive fluxes, whereas advection is negligible in the stationary state (Fig. S6).
@@
We showed that a chemical conversion reaction can arrest coarsening at various length scales. Generally, stronger conversion (larger $k$) leads to smaller length scales, although a range of scales are stable for each value of $k$. The final length scale thus also depends on initial conditions, and particularly the number of droplets that are present. We find that this initial droplet count depends strongly on advection, whereas advection is negligible in later stages. However, recent work suggests that even stronger reactions and advection cause a dynamic instability leading to spatiotemporal chaos Datt et al. (2025).
@@
Our work shows how regular patterns of droplets can be controlled by an interplay of chemical conversion and advection. Since these processes are crucial for biomolecular condensates, we speculate that cells exploit the discussed mechanism to control their condensates. However, cellular systems are much more complex, and it is thus likely that additional processes affect condensates Zwicker et al. (2025). In particular, cellular flows, e.g., driven by molecular motors, could further affect coarsening Stansell et al. (2006); Stratford et al. (2007); Gubbala et al. (2024), and inertia could be important on larger length scales Furtado (2006); Fielding (2008). Moreover, cellular droplets exhibit significant thermal fluctuations, which would not only be relevant for nucleation Ziethen et al. (2023), but could particularly allow the system to transition from one metastable state to another. In this case, we hypothesize that the most likely state exhibits a length scale close to $L_{\mathrm{eq}}$ given by Eq. (6). The multistability should then be interpreted as kinetic traps, potentially leading to glassy dynamics Zhang (2006).
@@
Acknowledgements—We thank Noah Ziethen and Riccardo Rossetto for helpful discussions. We gratefully acknowledge funding from the Max Planck Society and the European Research Council (ERC, EmulSim, 101044662). S. K. acknowledges funding through a fellowship of the IMPRS for Physics of Biological and Complex Systems. S. K. acknowledges support by the study program “Biological Physics” of the Elite Network of Bavaria.
@@ Appendix A Dimensionless equations and numerical details
To solve the dynamics of $c$ given by Eq. (7a) numerically, we use a finite difference method implemented in the py-pde Python package Zwicker (2020). During each time step, we use the projection method presented above using a pseudo-spectral method to determine $\bm{v}$. We use a Cartesian grid of size $(128\times 128)\sqrt{3}w$ for $k=0$ and $k=10^{-3}/\tau$, and $(256\times 256)\sqrt{3}w$ for $k=10^{-4}/\tau$ with resolution $\Delta x=\sqrt{3}w$. In our simulation we choose $w=\frac{1}{\sqrt{3}}$, $\tau=\frac{1}{9}$, an initial time step $\Delta t=0.045\tau$, and we use adaptive time stepping. We initialize the system with an average concentration $\bar{c}_{\mathrm{init}}=c_{0}$ with added noise drawn from a uniform distribution. For each Péclet number $\mathrm{P\hskip-1.0pte}$ we run 20 simulations to report means and standard deviations.
@@ A.1 Effective simulations
We use an effective droplet simulation method Kulkarni et al. (2023) to run simulations with low reaction rates, where the system volume needs to be large to get proper statistics, and the simulation time increases with $k^{-1}$,
@@ Appendix E Dissipation rates and predicted scaling
Figure S6: Velocity field of a hexagonal unit cell for $k=10^{-3}/\tau$, $c_{0}=0.4\,c_{\mathrm{in}}$ and $\mathrm{P\hskip-1.0pte}=100$. The arrow in the lower right corner shows the arrow-length of the maximum velocity. The viscous dissipation per unit volume is around 4 orders of magnitude smaller than the diffusive dissipation per unit volume (exact ratio after $T=900\tau$ is $q_{\mathrm{visc}}/q_{\mathrm{diff}}=2.30\cdot 10^{-4}$).

### v2

@@ II.3 Hydrodynamic advection accelerates coarsening by inducing coalescence
The relative importance of advection and diffusion can be quantified by the viscous and diffusive dissipation, respectively. Fig. 4D–F show that advection is relevant early, but decays with time, essentially independent of $k$ (blue lines are similar in panels D–F). The diffusive dissipation also peaks early, and decays similarly to the viscous dissipation during droplet coarsening. Interestingly, both dissipations are similar during the time where we observe a plateau in the pattern size (Fig. 4A–C). Beyond this point, diffusive dissipation dominates, and it plateaus at a much larger value compared to the viscous dissipation. This behavior indicates that energy induced by reactions leads to diffusive fluxes, whereas advection is negligible in the stationary state. This might be surprising since diffusive fluxes and the velocity field are driven by gradients in chemical potential. However, the fluid is incompressible ($\nabla\cdot\bm{v}=0$), so that chemical potential gradients get projected onto a solenoidal subspace. Therefore, osmotic pressure gradients are balanced by gradients in hydrostatic pressure. Accordingly, the velocity field remains small compared to diffusive fluxes for moderate advection strength (Fig. S6). In contrast, large advection (large $\mathrm{P\hskip-1.0pte}$) causes a dynamic instability leading to spatiotemporal chaos Datt et al. (2025).
III Discussion
@@
We showed that a chemical conversion reaction can arrest coarsening at various length scales. Generally, stronger conversion (larger $k$) leads to smaller length scales, although a range of scales are stable for each value of $k$. The final length scale thus also depends on initial conditions, and particularly the number of droplets that are present. We find that this initial droplet count depends strongly on advection, whereas advection is negligible in later stages.
@@
Our work shows how regular patterns of droplets can be controlled by an interplay of chemical conversion and advection. These processes are directly relevant for synthetic systems of chemically active droplets Jambon-Puillet et al. (2024); Sastre et al. (2024), and they might affect biomolecular condensates in cells. We thus speculate that cells exploit the discussed mechanism to control their condensates. However, cellular systems are much more complex, and it is thus likely that additional processes affect condensates Zwicker et al. (2025). In particular, cellular flows, e.g., driven by molecular motors, could further affect coarsening Stansell et al. (2006); Stratford et al. (2007); Gubbala et al. (2024), and inertia could be important on larger length scales Furtado (2006); Fielding (2008). Moreover, cellular droplets exhibit significant thermal fluctuations, which would not only be relevant for nucleation Ziethen et al. (2023), but could particularly allow the system to transition from one metastable state to another. In this case, we hypothesize that the most likely state exhibits a length scale close to $L_{\mathrm{eq}}$ given by Eq. (6). The multistability should then be interpreted as kinetic traps, potentially leading to glassy dynamics Zhang (2006).
Acknowledgements
@@
We thank Noah Ziethen and Riccardo Rossetto for helpful discussions. We gratefully acknowledge funding from the Max Planck Society and the European Research Council (ERC, EmulSim, 101044662). S. K. acknowledges funding through a fellowship of the IMPRS for Physics of Biological and Complex Systems. S. K. acknowledges support by the study program “Biological Physics” of the Elitenetzwerk Bayern.
@@ Appendix A Dimensionless equations and numerical details
Eqs. (7a)-(7b) are solved in a two-dimensional periodic domain. To solve the dynamics of $c$ given by Eq. (7a) numerically, we use a finite difference method implemented in the py-pde Python package Zwicker (2020). During each time step, we use the projection method presented above using a pseudo-spectral method to determine $\bm{v}$. We use a Cartesian grid of size $(128\times 128)\sqrt{3}w$ for $k=0$ and $k=10^{-3}/\tau$, and $(256\times 256)\sqrt{3}w$ for $k=10^{-4}/\tau$ with resolution $\Delta x=\sqrt{3}w$. In our simulation we choose $w=\frac{1}{\sqrt{3}}$, $\tau=\frac{1}{9}$, an initial time step $\Delta t=0.045\tau$, and we use adaptive time stepping. We initialize the system with an average concentration $\bar{c}_{\mathrm{init}}=c_{0}$ with added noise drawn from a uniform distribution. For each Péclet number $\mathrm{P\hskip-1.0pte}$ we run 20 simulations to report means and standard deviations.
@@ A.1 Effective simulations
We use an effective droplet simulation method Kulkarni et al. (2023) to run simulations in 2D with low reaction rates, where the system volume needs to be large to get proper statistics, and the simulation time increases with $k^{-1}$,
@@ Appendix E Dissipation rates and predicted scaling
Figure S6: (A) Hydrodynamic velocity field $\bm{v}$. (B) Diffusive flux velocity field $\bm{j}_{\mathrm{diff}}/c$. (A–B) Both snapshots show the stationary state in a hexagonal unit cell for $k=10^{-3}/\tau$, $c_{0}=0.4\,c_{\mathrm{in}}$, and $\mathrm{P\hskip-1.0pte}=100$. The legend in the lower right corner of each panel indicates that diffusive velocities are much larger, explaining why viscous dissipation is around 4 orders of magnitude smaller than the diffusive dissipation (ratio after $T=900\tau$ is $q_{\mathrm{visc}}/q_{\mathrm{diff}}=2.30\cdot 10^{-4}$).